OpenStudy (anonymous):

consider the equation C=5/9F-100/9. (C-Celsius, F-Fahrenheit) a) find the slope and C-intercept of this equation. b) the equation for F and convert 37 degree C into F-scale.

OpenStudy (saifoo.khan):

not a C intercpt, its y-intercpt. that is -100/9

OpenStudy (saifoo.khan):

Slope = 5/9

OpenStudy (saifoo.khan):

b) C=5/9F-100/9 37 + 100/9 = 5/9F F = 433/5 F = 86.6*F

OpenStudy (anonymous):

thanks alot man! actually im auditing and taxation student. thats why im asking things from you guys..

OpenStudy (saifoo.khan):

OIC, u from?

OpenStudy (saifoo.khan):

thats a Np man!

OpenStudy (anonymous):

Im from Islamabad, Pakistan.

OpenStudy (saifoo.khan):

Khi, Pakistan!